# Modeling Temporal Dependencies within the Target for Long-Term Time Series Forecasting

### v2

@@ Abstract
Long-term time series forecasting (LTSF) is a critical task across diverse domains. Despite significant advancements in LTSF research, we identify a performance bottleneck in existing LTSF methods caused by the inadequate modeling of Temporal Dependencies within the Target (TDT). To address this issue, we propose a novel and generic temporal modeling framework, Temporal Dependency Alignment (TDAlign), that equips existing LTSF methods with TDT learning capabilities. TDAlign introduces two key innovations: 1) a loss function that aligns the change values between adjacent time steps in the predictions with those in the target, ensuring consistency with variation patterns, and 2) an adaptive loss balancing strategy that seamlessly integrates the new loss function with existing LTSF methods without introducing additional learnable parameters. As a plug-and-play framework, TDAlign enhances existing methods with minimal computational overhead, featuring only linear time complexity and constant space complexity relative to the prediction length. Extensive experiments on six strong LTSF baselines across seven real-world datasets demonstrate the effectiveness and flexibility of TDAlign. On average, TDAlign reduces baseline prediction errors by 1.47% to 9.19% and change value errors by 4.57% to 15.78%, highlighting its substantial performance improvements.
Index Terms:
Time series analysis, long-term forecasting, temporal dependency.
@@ I Introduction
In this work, we propose Temporal Dependency Alignment (TDAlign), a generic framework that enables existing methods to capture temporal dependencies within the target time series via an additional learning objective. We first formulate the TDT with the change values between adjacent time steps to facilitate capturing intrinsic temporal patterns of the target. Building upon this, we introduce a novel loss function to align the change values within the prediction with those within the target time series, thereby guiding the model to effectively learn intricate temporal relations. Finally, we design an adaptive loss balancing strategy that dynamically adjusts the relative importance between the conventional forecasting objective and the TDT learning objective, enabling TDAlign to flexibly integrate with existing LTSF methods without introducing additional learnable parameters (Section III-C3). Computational complexity analysis demonstrates that TDAlign incurs negligible computational overhead, with time complexity of $\mathcal{O}(H)$ and space complexity of $\mathcal{O}(1)$, where $H$ represents the length of the prediction (Section III-D).
Flexibility and Effectiveness. Remarkably, our proposed TDAlign is orthogonal to existing LTSF methods without introducing additional learnable parameters. We evaluate TDAlign using a broad spectrum of strong baseline methods, including the MLP-based method DLinear [14], CNN-based methods MICN [17] and TimesNet [18], Transformer-based methods PatchTST [20] and iTransformer [29], and the RNN-based method SegRNN [22]. Experimental results on seven widely used LTSF benchmark datasets demonstrate that TDAlign consistently improves the forecasting performance of these six baseline methods, showcasing its strong flexibility and effectiveness (Section IV-B). As shown in Figure 1, TDAlign achieves significant error reductions across seven datasets: (1) For change values between adjacent time steps within the prediction, absolute error reductions range from 4.57% to 15.78%. (2) For the prediction, absolute error reductions range from 1.47% to 9.19%.
@@ III-A Preliminaries
In the following, input refers to the historical time series, while target and prediction represent the ground truth and predicted values of the future time series, respectively. Let $x_{t}\in\mathbb{R}^{N}$ denote the observation at time step $t$ of $N$ variables. In LTSF tasks, given the input $X=\left\{x_{t-L+1},x_{t-L+2},\cdots,x_{t}\right\}\in\mathbb{R}^{L\times N}$, the model $\mathcal{F}$ with learnable parameters $\theta$ aims to predict the target $Y=\{x_{t+1},x_{t+2},\dots,x_{t+H}\}$, where the corresponding prediction is $\widehat{Y}=\left\{\hat{x}_{t+1},\hat{x}_{t+2},\cdots,\hat{x}_{t+H}\right\}\in\mathbb{R}^{H\times N}$. Here, $L$ is the length of the input, while $H$ is the length of the target and the prediction. Typically, $H$ is much larger in LTSF tasks than in traditional short-term time series forecasting tasks, such as 720 time steps. This forecast behavior can be formally summarized as $\mathcal{F}_{\theta}:X\rightarrow\widehat{Y}$. Our ultimate goal is to train a model that minimizes the discrepancy between the prediction $\widehat{Y}$ and the target $Y$. Frequently used notations are listed in Table I.
@@ III-B Motivation
An intrinsic feature of time series is that consecutive time steps typically exhibit strong correlations [44]. This inherent characteristic suggests that forecasting models should adequately account for temporal dependencies across multiple steps, encompassing both the input and target [45]. Inspired by this insight, we assess the effort of existing methods in capturing temporal variations.
@@
$\mathcal{L}_{Y}=\frac{1}{H}\sum^{H}_{i=1}\ell(x_{t+i},\hat{x}_{t+i})$ (1)
where $\ell(\cdot,\cdot)$ denotes the error evaluated at a single time step. Popular choices for $\mathcal{L}_{Y}$ include $L1$ loss, $L2$ loss and other variants [48]. However, methods merely optimize the model towards the minimum of the average of step-wise prediction error with $\mathcal{L}_{Y}$, overlooking the temporal relations when generating prediction steps [49, 50].
@@ III-C1 Temporal Dependencies within the Target
A fair formulation of temporal dependencies should benefit not only the learning of these dependencies but also the evaluation used to assess the learning effects. We consider two key insights: (1) In time series, temporally proximate time steps typically exhibit stronger correlations compared to those with far temporal distance [44, 51, 45]; (2) The autoregressive decoding strategy enables methods to naturally learn sequential dependencies, where each prediction step is generated conditioned on its immediately preceding prediction step [25, 26].
Inspired by these insights, we formulate Temporal Dependencies within the Target (TDT) using first-order differencing values [52] of the target. Specifically, we formulate TDT as $D=\left\{d_{t+1},d_{t+2},\cdots,d_{t+H}\right\}\in\mathbb{R}^{H\times N}$, where each term $d_{t+i}\in\mathbb{R}^{N}$ equals to the change value between adjacent time steps, calculated as:
@@
Despite its simplicity, using first-order differencing values (i.e., change values between adjacent time steps) to formulate temporal dependencies demonstrates greater generality and effectiveness compared to variants that use higher-order differencing values or change values between non-adjacent time steps (see the experiments in Section IV-C for details).
@@ III-C3 Adaptive Loss Balancing
To enable TDAlign to integrate with the forecasting model, an intuitive solution is to set a weight $\lambda$ for balancing the learning procedure, which can be expressed as:
$\mathcal{L}=\lambda\cdot\mathcal{L}_{Y}+(1-\lambda)\cdot\mathcal{L}_{D}.$ (5)
While manually tuning $\lambda$ to determine its optimal value is a straightforward option, this approach lacks flexibility and is labor-intensive. Therefore, we propose an adaptive loss balancing strategy using a dynamic weight $\rho$, which is calculated as follows:
$\rho=\frac{1}{H}\sum_{i=1}^{H}\mathbbm{1}(\operatorname{sgn}(d_{t+i})\neq\operatorname{sgn}(\hat{d}_{t+i}))$ (6)
@@
Figure 3: Overview of our proposed Temporal Dependency Alignment (TDAlign) framework. First, TDAlign formulates temporal dependencies within the target (TDT)—denoted as $D$, and temporal dependencies within the prediction (TDP)—denoted as $\widehat{D}$, using the first-order differencing values of the target $Y$ and the prediction $\widehat{Y}$, respectively (in green boxes). Then, TDAlign devises a loss $L_{D}$ that guides the model to align $\widehat{D}$ with $D$ for capturing temporal patterns. Finally, TDAlign calculates a balance weight $\rho$ (in the orange box) to adaptively control the relative contributions of the conventional forecasting loss $L_{Y}$ and the devised $L_{D}$, thereby improving existing methods without introducing additional learnable parameters.
The dynamic weight $\rho$ offers several advantages: (1) Simplicity and informativeness. $\rho$ is directly computed from TDP and TDT without requiring additional data or complex preprocessing. Moreover, the sign indicates the change direction between adjacent time steps. Therefore, $\rho$ can also serve as a metric to quantify the method’s ability to capture TDT, as detailed in Section IV-A3. (2) Effectiveness. $\rho$ dynamically adjusts the relative importance between the conventional forecasting objective $\mathcal{L}_{Y}$ (Equation 1) and the TDT learning objective $\mathcal{L}_{D}$ (Equation 4) based on the model’s current learning state. The effectiveness of $\rho$ is further validated by experiments detailed in Section IV-C. (3) Flexibility. The formulation of $\rho$ introduces no additional learnable parameters, enabling TDAlign to serve as a model-agnostic framework that seamlessly integrates with various LTSF methods.
III-C4 Training
Formally, when integrating with existing LTSF methods, the overall loss function of TDAlign is expressed as:
$\mathcal{L}=\rho\cdot\mathcal{L}_{Y}+(1-\rho)\cdot\mathcal{L}_{D}.$ (7)
@@ III-C4 Training
       Compute the overall loss $\mathcal{L}\leftarrow\rho\cdot\mathcal{L}_{Y}+(1-\rho)\cdot\mathcal{L}_{D}$;
@@
Algorithm 1 Temporal Dependencies Alignment
@@ III-D Computational Complexity Analysis
In conclusion, TDAlign is an efficient plug-and-play framework, introducing minimal computational overhead that scales linearly with prediction length in time complexity, while maintaining constant space complexity.
IV Experiments
@@
In this section, we conduct extensive experiments to demonstrate the effectiveness and generalizability of our proposed TDAlign. We begin with a detailed description of the experimental setup. Subsequently, we present the main experimental results and statistics of improvements achieved by integrating TDAlign into baseline methods. Furthermore, we conduct ablation studies to evaluate the effectiveness of various components of TDAlign and assess the impact of the formulation of temporal dependencies. Finally, we showcase several long-term forecasting instances to demonstrate that TDAlign produces more accurate and realistic predictions compared to baseline methods.
IV-A Experimental Setup
IV-A1 Datasets

### v3

@@ Abstract
Long-term time series forecasting (LTSF) is a critical task across diverse domains. Despite significant advancements in LTSF research, we identify a performance bottleneck in existing LTSF methods caused by the inadequate modeling of Temporal Dependencies within the Target (TDT). To address this issue, we propose a novel and generic temporal modeling framework, Temporal Dependency Alignment (TDAlign), that equips existing LTSF methods with TDT learning capabilities. TDAlign introduces two key innovations: 1) a loss function that aligns the change values between adjacent time steps in the predictions with those in the target, ensuring consistency with variation patterns, and 2) an adaptive loss balancing strategy that seamlessly integrates the new loss function with existing LTSF methods without introducing additional learnable parameters. As a plug-and-play framework, TDAlign enhances existing methods with minimal computational overhead, featuring only linear time complexity and constant space complexity relative to the prediction length. Extensive experiments on six strong LTSF baselines across seven real-world datasets demonstrate the effectiveness and flexibility of TDAlign. On average, TDAlign reduces baseline prediction errors by 1.47% to 9.19% and change value errors by 4.57% to 15.78%, highlighting its substantial performance improvements. We make our codes publicly available at https://github.com/XQ-edu/TDAlign.
@@ I Introduction
In this work, we propose Temporal Dependency Alignment (TDAlign), a generic framework that enables existing methods to capture temporal dependencies within the target time series via an additional learning objective. In Section III-C, we first formulate the TDT with the change values between adjacent time steps to facilitate capturing intrinsic temporal patterns of the target. Building upon this, we introduce a novel loss function to align the change values within the prediction with those within the target time series, thereby guiding the model to effectively learn intricate temporal relations. Finally, we design an adaptive loss balancing strategy that dynamically adjusts the relative importance between the conventional forecasting objective and the TDT learning objective, enabling TDAlign to flexibly integrate with existing LTSF methods without introducing additional learnable parameters (Section III-C3). We provide further theoretical analysis of TDAlign in Section IV. Computational complexity analysis in Section III-D demonstrates that TDAlign incurs negligible computational overhead, with time complexity of $\mathcal{O}(H)$ and space complexity of $\mathcal{O}(1)$, where $H$ represents the length of the prediction.
Flexibility and Effectiveness. Remarkably, our proposed TDAlign is orthogonal to existing LTSF methods without introducing additional learnable parameters. We evaluate TDAlign using a broad spectrum of strong baseline methods, including the MLP-based method DLinear [14], CNN-based methods MICN [17] and TimesNet [18], Transformer-based methods PatchTST [20] and iTransformer [29], and the RNN-based method SegRNN [22]. Experimental results on seven widely used LTSF benchmark datasets demonstrate that TDAlign consistently improves the forecasting performance of these six baseline methods, showcasing its strong flexibility and effectiveness (Section V-B). As shown in Figure 1, TDAlign achieves significant error reductions across seven datasets: (1) For change values between adjacent time steps within the prediction, absolute error reductions range from 4.57% to 15.78%. (2) For the prediction, absolute error reductions range from 1.47% to 9.19%.
@@ III-A Preliminaries
In the following, input refers to the historical time series, while target and prediction represent the ground truth and predicted values of the future time series, respectively. Let $x_{t}\in\mathbb{R}^{N}$ denote the observation at time step $t$ of $N$ variables. In LTSF tasks, given the input $X=\left\{x_{t-L+1},x_{t-L+2},\cdots,x_{t}\right\}\in\mathbb{R}^{L\times N}$, the model $\mathcal{F}$ with learnable parameters $\theta$ aims to predict the target $Y=\{x_{t+1},x_{t+2},\dots,x_{t+H}\}$, where the corresponding prediction is $\widehat{Y}=\left\{\hat{x}_{t+1},\hat{x}_{t+2},\cdots,\hat{x}_{t+H}\right\}\in\mathbb{R}^{H\times N}$. Here, $L$ is the length of the input, while $H$ is the length of the target and the prediction. Typically, $H$ is much larger in LTSF tasks than in traditional short-term time series forecasting tasks, such as 720 time steps. This forecast behavior can be formally summarized as $\mathcal{F}_{\theta}:X\rightarrow\widehat{Y}$.
Our ultimate goal is to train a model that minimizes the discrepancy between the prediction $\widehat{Y}$ and the target $Y$. Frequently used notations are listed in Table I.
@@ III-B Motivation
Minimizing the discrepancy between the prediction and the target requires not only achieving point-wise accuracy but also capturing the temporal dynamics inherent in the target [43].
@@
$\mathcal{L}_{Y}=\frac{1}{H}\sum^{H}_{i=1}\ell(x_{t+i},\hat{x}_{t+i}),$ (1)
where $\ell(\cdot,\cdot)$ denotes the error evaluated at a single time step. Popular choices for $\mathcal{L}_{Y}$ include MSE, MAE and other variants [46]. However, methods merely optimize the model towards the minimum of the average of point-wise prediction error with $\mathcal{L}_{Y}$, overlooking the temporal relations when generating prediction steps [47, 48].
@@ III-C1 Temporal Dependencies within the Target
A fair formulation of temporal dependencies should benefit both learnability and generalizability in capturing these dependencies. To achieve this, we consider two key factors. First, in time series, temporally proximate time steps typically exhibit stronger correlations compared to those with far temporal distance [49, 50, 43]. Consequently, change patterns between adjacent time steps contain more direct and reliable information than those across longer intervals. Second, while some formulations could capture specific dependencies, such as using higher-order differences for complex relations, long-range intervals for periodic patterns [18], or frequency-domain features for global structures [51], they face inherent limitations. Specifically, they often incorporate strong assumptions tailored to specific data characteristics, limiting their transferability across different domains. Moreover, their increased complexity makes them more susceptible to overfitting spurious patterns or amplifying noise, which can be detrimental to the baseline model’s performance.
Building upon these considerations, we formulate Temporal Dependencies within the Target (TDT) using first-order differencing values [52] of the target. This formulation captures essential local temporal dynamics while maintaining simplicity and broad applicability without imposing restrictive data-specific assumptions. Specifically, we formulate TDT as $D=\left\{d_{t+1},d_{t+2},\cdots,d_{t+H}\right\}\in\mathbb{R}^{H\times N}$, where each term $d_{t+i}\in\mathbb{R}^{N}$ equals to the change value between adjacent time steps, calculated as:
@@
Despite its simplicity, this formulation using first-order differencing values (i.e., change values between adjacent time steps) demonstrates greater generality and effectiveness compared to alternative formulations of temporal dependencies, such as higher-order differencing values, change values between non-adjacent time steps, or frequency domain features (see the experiments in Section V-C for details).
@@ III-C3 Adaptive Loss Balancing
To integrate TDAlign with forecasting models, a prevalent strategy is to introduce a weight $\rho$ to balance the relative importance between the conventional forecasting objective $\mathcal{L}_{Y}$ and the TDT learning objective $\mathcal{L}_{D}$. Thus, the overall loss function of TDAlign can be expressed as:
$\mathcal{L}=\rho\mathcal{L}_{Y}+(1-\rho)\mathcal{L}_{D}.$ (5)
While manually tuning $\rho$ is a straightforward option, this process lacks flexibility and is labor-intensive. Alternatively, treating $\rho$ as a trainable parameter offers a more automated solution. However, it cannot adaptively modulate the relative contributions in response to evolving temporal dynamics within the data or the model’s varying performance across different training phases. To address these limitations, we propose an adaptive loss balancing strategy using a dynamic weight $\rho$, which is updated based on the model’s instantaneous performance on the TDT learning task and calculated as follows:
$\rho=\frac{1}{H}\sum_{i=1}^{H}\mathbbm{1}(\operatorname{sgn}(d_{t+i})\neq\operatorname{sgn}(\hat{d}_{t+i})),$ (6)
@@
Figure 3: Overview of our proposed Temporal Dependency Alignment (TDAlign) framework. First, TDAlign formulates temporal dependencies within the target (TDT)—denoted as $D$, and temporal dependencies within the prediction (TDP)—denoted as $\widehat{D}$, using the first-order differencing values of the target $Y$ and the prediction $\widehat{Y}$, respectively (in green boxes). Then, TDAlign devises a loss $\mathcal{L}_{D}$ that guides the model to align $\widehat{D}$ with $D$ for capturing temporal patterns. Finally, TDAlign calculates a balance weight $\rho$ (in the orange box) to adaptively control the relative contributions of the conventional forecasting loss $\mathcal{L}_{Y}$ and the devised $\mathcal{L}_{D}$, thereby improving existing methods without introducing additional learnable parameters.
The dynamic weight $\rho$ offers several advantages: (1) Simplicity and informativeness. $\rho$ is directly computed from TDP and TDT without requiring additional data or complex preprocessing. Moreover, the sign indicates the change direction between adjacent time steps. Therefore, $\rho$ can also serve as a metric to quantify the method’s ability to capture TDT, as detailed in Section V-A3. (2) Effectiveness. $\rho$ dynamically adjusts the relative importance between the conventional forecasting objective $\mathcal{L}_{Y}$ and the TDT learning objective $\mathcal{L}_{D}$ based on the model’s current learning state. The effectiveness of $\rho$ is further theoretically analyzed in Section IV and validated by experiments detailed in Section V-C. (3) Flexibility. The formulation of $\rho$ introduces no additional learnable parameters, enabling TDAlign to serve as a model-agnostic framework that seamlessly integrates with various LTSF methods.
@@
   Compute the overall loss $\mathcal{L}\leftarrow\rho\mathcal{L}_{Y}+(1-\rho)\mathcal{L}_{D}$;
@@
Algorithm 1 Temporal Dependency Alignment
@@ III-D Computational Complexity Analysis
In conclusion, TDAlign is an efficient plug-and-play framework, introducing minimal computational overhead that scales linearly with prediction length in time complexity, while maintaining constant space complexity. We empirically validate these efficiency claims in Section V-C3.
IV Theoretical Analysis
For computational efficiency, conventional LTSF methods adopt non-autoregressive decoding, generating all prediction steps in a single forward pass. They assume temporal independence across prediction steps and optimize solely with a point-wise loss function $\mathcal{L}_{Y}$. However, such independence assumptions fundamentally neglect the temporal correlations inherent in time series, leading to a critical paradox: while these methods may achieve low numerical errors like MSE, they often generate predictions with unrealistic high-frequency fluctuations. This behavior contradicts the temporal coherence of real-world time series, which typically evolve through smooth, locally consistent transitions rather than abrupt, erratic variations.
To formalize this limitation, we quantify temporal dependencies that $\mathcal{L}_{Y}$ fails to capture. Following the maximum likelihood principle, we analyze the discrepancy $\Psi$ between $\mathcal{L}_{Y}$ and the Negative Log-Likelihood (NLL) objective under a first-order Markov assumption. This assumption is well-justified as it directly models the underlying generative process of many time series, where the current state predominantly determines the next [49]. This fundamental principle not only manifests as strong dependencies between adjacent time steps but also forms the foundational basis of autoregressive modeling.
Theorem 1. Under a first-order Markov assumption, where $p(x_{t+i}|x_{<t+i},X)\approx p(x_{t+i}|x_{t+i-1},X)$, the discrepancy $\Psi$ between the point-wise MSE loss, i.e., $\mathcal{L}_{Y}$, and the NLL objective is given by: $\Psi=\sum_{i=2}^{H}\frac{1}{1-\phi_{i}^{2}}\left[\phi_{i}^{2}(\epsilon_{i}^{2}+\epsilon_{i-1}^{2})-2\phi_{i}\epsilon_{i}\epsilon_{i-1}\right],$ (7) where $\epsilon_{i}=x_{t+i}-\hat{x}_{t+i}$ is the prediction error at step $i$, and $\phi_{i}$ is the partial autocorrelation coefficient between $x_{t+i}$ and $x_{t+i-1}$.
A detailed proof is provided in Appendix. Theorem 1 reveals a key limitation of the conventional objective $\mathcal{L}_{Y}$: failing to account for the temporal dependencies among time steps, mathematically quantified by the discrepancy term $\Psi$. This omission explains why models trained solely on $\mathcal{L}_{Y}$ tend to produce predictions with insufficient temporal coherence. However, directly optimizing $\Psi$ is practically infeasible. The primary barriers are the prohibitive computational cost and numerical instability associated with estimating the partial autocorrelation coefficients, $\phi_{i}$, especially for non-stationary time series [49, 50].
Our proposed loss, $\mathcal{L}_{D}$, is a tractable surrogate objective that directly addresses the above limitation. Specifically, when implemented using MSE, $\mathcal{L}_{D}=\sum(d_{t+i}-\hat{d}_{t+i})^{2}=\sum((x_{t+i}-x_{t+i-1})-(\hat{x}_{t+i}-\hat{x}_{t+i-1}))^{2}=\sum(\epsilon_{i}-\epsilon_{i-1})^{2}$. By penalizing the difference between consecutive errors, $\mathcal{L}_{D}$ mirrors the core principle of $\Psi$ without its computational complexity. Therefore, $\mathcal{L}_{D}$ explicitly constrains the model to capture step-wise temporal dynamics.
To effectively integrate $\mathcal{L}_{Y}$ and $\mathcal{L}_{D}$, we propose an adaptive loss balancing strategy using $\rho$, which modulates the contribution of each loss term according to the model’s performance on the TDT learning task during training.
Proposition 1. Let the prediction error $e_{i}=d_{t+i}-\hat{d}_{i+1}$ as a zero-mean Gaussian variable with variance $\sigma_{e}^{2}$, the expected value of the adaptive weight $\rho$ is given by: $\mathbb{E}[\rho|\sigma_{e}^{2}]=\frac{1}{H}\sum_{i=1}^{H}\Phi\left(-\frac{|d_{t+i}|}{\sigma_{e}}\right),$ (8) where $\Phi(\cdot)$ is the cumulative distribution function of the standard normal distribution, $|d_{t+i}|$ is the true change magnitude between $x_{t+i}$ and $x_{t+i-1}$, and $\sigma_{e}$ is the standard deviation of the model’s prediction error.
A detailed proof is provided in Appendix. Proposition 1 reveals a direct relationship between the expected dynamic weight, $\mathbb{E}[\rho]$, and the error variance, $\sigma_{e}^{2}$, which serves as a proxy for the model’s performance on the TDT learning task. This relationship dictates a dynamic balancing of the loss terms throughout training. Initially, when the model’s predictions are inaccurate, the large error variance $\sigma_{e}^{2}$ leads to a higher $\mathbb{E}[\rho]$. This prioritizes $\mathcal{L}_{Y}$, thereby guiding the model to capture the global pattern of the target sequence. As training progresses, the model’s predictive accuracy improves, causing $\sigma_{e}^{2}$ and consequently $\mathbb{E}[\rho]$ to decrease. This reduction gradually shifts the optimization focus towards $\mathcal{L}_{D}$, encouraging the refinement of predictions by capturing fine-grained local dynamics. In essence, $\rho$ dynamically steers the training focus from learning coarse, global patterns toward capturing fine-grained, local details. This adaptive loss balancing facilitates the model to build richer sequence representations, thereby enhancing its generalization capabilities.
V Experiments
@@
In this section, we conduct extensive experiments to demonstrate the effectiveness and generalizability of our proposed TDAlign. We begin with a detailed description of the experimental setup. Subsequently, we present the main experimental results and statistics of improvements achieved by integrating TDAlign into baseline methods. Furthermore, we conduct ablation studies to evaluate the effectiveness of various components of TDAlign and assess the impact of the formulation of temporal dependencies. Additionally, we perform additional investigations into the generalizability and robustness of TDAlign. Finally, we showcase several long-term forecasting instances to demonstrate that TDAlign produces more accurate and realistic predictions compared to baseline methods.
V-A Experimental Setup
V-A1 Datasets